OpenStudy (anonymous):

Martin rides to Cambridge 30 miles away from home at the average rate of 10 miles per hour. He return on the better road that is 50% longer where he can increase his rate by 100%. How much time does he save by taking the better road on his return trip? a-2 1/4 mins b-15 mins c-45 mins d-2 hrs 15 mins e-2 hrs 45 mins I suck at these kind of questions, help?

OpenStudy (phi):

30 miles at 10 miles per hour if you use rate * time = distance can you figure out how long it took to get to Cambridge ?

OpenStudy (anonymous):

10?

OpenStudy (phi):

replace rae with 10, time with T (unknown), and distance with 30 10 T = 30 to solve for T (in hours) divide both sides by 10 can you do that ?

OpenStudy (anonymous):

Wait, no 3 hrs?

OpenStudy (phi):

yes, 3 hours now part two, returning the road that is 50% longer than 30 miles what is 50% of 30 ?

OpenStudy (anonymous):

15.

OpenStudy (phi):

so the road back is 15 longer than 30 how long is the road back ?

OpenStudy (anonymous):

45...?

OpenStudy (phi):

yes, so distance back is 45 increase his rate by 100% his rate was 10 mph. if you increase by 100% you increase by 10 his rate going home is 10+10= 20 mph rate * time= distance 20 * T= 45 how long does it take to get home (in hours) ?

OpenStudy (phi):

20 * T= 45 to solve, divide both sides by 20

OpenStudy (anonymous):

2 hrs and 15 minutes?

OpenStudy (phi):

yes, 2 and a quarter hours He took 3 hours to get there and 2 hours 15 mins to get back How much time does he save by taking the better road on his return trip?

OpenStudy (anonymous):

45 mins